Accelerator Update for October 31, 2008 - November 3, 2008
Friday Oct 31
The day shift began with the Operations monitoring the Tevatron (TeV) store 6533 (252.55E30) and the Recycler stash (54.2E10), with the Antiproton Source (Pbar) stacking (19.7E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE, NuMI, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
At 8:14 AM, Operations reported that no beam was making it into the AP2 transfer line. Pbar experts said that this was due to a target problem.
Linac experts began a Linac RF station (LRF5) study at 10:01 AM.
A TeV expert reported at 10:32 AM that he had reset the antiprotons abort kicker. He said that he thought there was a poor connection. He will investigate this later.
At 3:08 PM, Pbar experts reported that the bolts holding the target sheered off. They began working on its replacement. Later, they will study the cause of the failure.
Operations disabled beam to MTest at 6:09 PM.
A Pbar expert reported at 7:01 PM that he had successfully pulse tested the target and technicians were replacing the target shield blocks.
Operations established beam to Pbar at 7:40 PM. A Pbar expert was conducting a target scan. He finished the scan about an hour later.
At 11:04 PM, a MI RF station (MIRF3) tripped off for the first of four times and then, after it would not reset, operators bypassed the station.
Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift.
Saturday Nov 1
The midnight shift began with Operations monitoring store 6533 (64.46E30) and the stash (147.7E10), with Pbar stacking (22.1E10), and with the MiniBooNE and NuMI experiments taking beam.
Operations terminated store 6533 at 5:11 AM.
Operations put the TeV into shot setup at 5:22 AM.
Operations established store 6534 at 6:43 AM with an initial luminosity of 256.99E30.
Operations established beam to MTest at 7:49 AM.
Operations reported at 8:02 AM that the Recycler has stashed seven times during the midnight shift.
At 8:42 AM, LRF3 tripped off for the first of two times in about an hour. It reset.
At 1:01 PM, Operations reported that the Pelletron VME and AccelNet came back, but the blower power supply would not come on. Experts investigated and brought ECOOL back online.
An RF expert reported at 4:51 PM that he had brought MIRF3 back online.
Operations reported at 8:20 PM that the Recycler had stashed nine times during the last 12 hours.
Operations terminated store 6534 at 10:55 PM.
Operations put the TeV into shot setup at 10:56 PM.
At 11:07 PM, Operations reported problems with beam positions that were causing MI beam aborts.
Sunday Nov 2
The midnight shift began with TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with the MiniBooNE and NuMI experiments taking beam.
Operations established store 6535 at 12:38 AM with an initial luminosity of 318.98E30. This store set a New Luminosity Record. The old record was 318.49E30.
Operations reported at 7:57 AM that the Recycler had stashed 10 times during the last 12 hours.
At 10:50 AM, Operations reported that the Linac beam intensity had dropped. Two Linac experts were called in.
Operations terminated store 6535 at 11:08 AM.
Operations prepared CDF and D0 for access at 11:38 AM.
At 12:39 PM, Operators reported loud banging coming from a Linac Klystron RF station (KRF3). Operations called in experts.
D0 and CDF came out of access at 3:17 PM.
Operations put the TeV through a dry squeeze at 4:09 PM.
At 4:22 PM, a Linac expert reported that he had fixed one problem with KRF3, but another problem turned up. Repairs continued.
Operations reported at 8 PM that the Recycler had stashed four times during the last 12 hours.
Linac experts reported at 9:02 PM that they had completed their repairs of KRF3. After completing several repairs they found a bad cable.
Operators completed a TeV orbit check at 9:57 PM.
Operations put the TeV through the first of two wet squeezes at 10:22 PM.
Operations put the TeV into shot setup at 11:17 PM.
Monday Nov 3
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with the MiniBooNE and NuMI experiments taking beam.
Operations established store 6539 at 12:47 AM with an initial luminosity of 277.47E30.
At 12:55 AM, Operations reported a problem with the flying wires that might be related to the time change.
At 6:42 AM, MIRF3 tripped off and would not reset. Operators bypassed the station. An RF expert came in at 7:57 AM and brought the station back online.
Operations reported at 8:01 AM that the Recycler had stashed 10 times during the last 12 hours.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 11/03/2008 email Fermilab|
|Security, Privacy, Legal|